# PRISM: Deriving the Transformer as a Signal-Denoising Operator via Maximum Coding Rate Reduction

### v1

@@ Abstract
Deep learning models, particularly Transformers, are often criticized as ”black boxes” and lack interpretability. We propose Prism, a white-box attention-based architecture derived from the principles of Maximizing Coding Rate Reduction ($\text{MCR}^{2}$). By modeling the attention mechanism as a gradient ascent process on a distinct signal-noise manifold, we introduce two physical constraints: an overcomplete dictionary to expand the representational phase space, and an irrational frequency separation ($\pi$-RoPE) to enforce incoherence between signal and noise subspaces. We demonstrate that these geometric inductive biases can be viewed as a physical constraint and they are sufficient to induce unsupervised functional disentanglement alone. Using TinyStories as a controlled testbed for verifying spectral dynamics, we observe that Prism spontaneously specializes its attention heads into spectrally distinct regimes: low-frequency heads capturing long-range causal dependencies (signal) and high-frequency heads handling local syntactic constraints (noise). Our results suggest that interpretability and performance are not a trade-off, but can be unified through principled geometric construction.
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) has established itself as the fundamental building block of modern artificial intelligence, exhibiting a remarkable power-law relationship between parameter scale and performance (Kaplan et al., 2020). However, despite their empirical success, these models operate largely as black boxes. One limitation is the architectural conflation between semantic reasoning (the ability to manipulate abstract concepts over long-range contexts) and syntactic memorization (the rote statistical completion of local patterns). While recent mechanistic interpretability efforts have identified specific circuits such as induction heads (Olsson et al., 2022), the standard attention mechanism lacks the physical inductive bias required to separate these distinct functional regimes by design. This may often result in brittle reasoning, hallucination driven by overfitting to noise and data inefficiency.
@@
To enhance interpretability, we turn to first-principles approaches that view deep learning as data compression. The principle of Maximal Coding Rate Reduction ($MCR^{2}$) (Chan et al., 2022) shows that the fundamental goal of representation learning is to expand the volume of the dataset globally while compressing the volume of similar classes locally. Building on this theoretical foundation, the CRATE (Coding-RATE transformer) framework (Yu et al., 2024) establishes that the self-attention mechanism is mathematically isomorphic to a step of gradient ascent optimizing the rate reduction. This perspective transforms the attention layer mechanism from an empirical heuristic into a derived operator for signal denoising on low-dimensional manifolds, providing the necessary white-box substrate for our architectural interventions.
@@
However, gaps between white-box neural networks and black-box neural networks still exist. One critical gap lies in data modeling; current white-box models typically assume clean data or Gaussian noise, whereas real-world data distributions are highly non-Gaussian. This may result in the coherence of optimization trajectory. If the model has limited capacity, the gradient flows of semantic signals (typically low-frequency global dependencies) and syntactic noise (high-frequency local statistics) mix together and hinder the learning. To mitigate this, we propose Prism, a physics-informed transformer architecture that introduces two specific geometric constraints to the standard Transformer block based on explicit additive noise. First, we impose an overcomplete dictionary (Candès et al., 2006) to expand the representational phase space, allowing signal and noise to separate in higher dimensions without competing for the same limited basis vectors. Second, we implement frequency separation via $\pi$-scaled Rotary Positional Embeddings ($\pi$-RoPE) (Su et al., 2024). By scaling the rotational bases of the overcomplete subspaces by $\pi$ and $1/\pi$ respectively, we hypothesize that this choice effectively mitigates resonance and keeps the reasoning and memorization dynamics on spectrally disjoint manifolds.
@@ 2.1 Rate Reduction as the Objective
and the learning objective is to extract $\mathbf{Z}_{0}$ from $\mathbf{Z}$. To achieve this, we follow the maximum coding rate reduction framework. Let $\mathbf{Z}\in\mathbb{R}^{N\times d}$ denote representation (e.g. input tokens), and $\mathbf{U}_{s},\mathbf{U}_{e}$ be the subspace with respect to the effective signal $\mathbf{Z}_{0}$ and noise $\mathbf{E}$ respectively. We assume the signal and noise reside in distinct subspaces, and these subspaces $\mathbf{U}_{s}$ and $\mathbf{U}_{n}$ are sufficiently incoherent (e.g., $\left\|\mathbf{U}_{s}^{\top}\mathbf{U}_{n}\right\|\approx 0$). We can approximate the noise $\mathbf{E}=\mathbf{U}_{n}^{\top}\mathbf{Z}$. Thus, the effective signal $Z_{0}$ is obtained by maximizing the coding rate difference $\Delta R(\mathbf{Z})$:
@@
where $R(\mathbf{Z})$ represents the coding rate of the signal projected onto local subspaces expanded by $\mathbf{U}_{s}$, and $R(\mathbf{Z}|\mathbf{U}_{n})$ represents the coding rate of the noise obtained by projecting the noisy signal onto the local subspaces $U_{n}$, $\lambda>0$ is the parameter controlling the separation between signal/noise decomposition. Specifically, under the Gaussian assumption, the coding rate of the matrix $\mathbf{Z}$ is defined as:
@@
If there are many subspaces $\mathbf{U}=[\mathbf{U}_{1},\dots,\mathbf{U}_{K}]$, the coding rate of $\mathbf{Z}$ is the sum of the coding rates of all subspace projections, and yields
$R(\mathbf{Z}\mathbf{U}_{k})=\frac{1}{2}\log\det\left(\mathbf{I}+\beta(\mathbf{Z}\mathbf{U}_{k})^{\top}(\mathbf{Z}\mathbf{U}_{k})\right),$ (4)
where $\beta=\frac{p}{N\epsilon^{2}}$ and $p$ is the dimension of subspaces.
@@ 2.2 Prism as Gradient Ascent of Coding Rate Reduction
Figure 1: The PRISM White-Box Architecture. (Top) The input latent state $\mathbf{Z}^{\ell}$ is projected into an overcomplete feature space via dictionary $\mathbf{U}=[\mathbf{U}_{s},\mathbf{U}_{n}]$. The data flow in prism bifurcates into Signal Stream (Blue) and Noise Stream (Orange). $\pi$-RoPE: the Signal stream applies low-frequency rotary embeddings scaled by $\pi$ ($e^{i\pi\theta}$) to represent long-range semantic structures, while the Noise stream applies high-frequency embeddings ($e^{i\pi^{-1}\theta}$) to capture short-range syntactic artifacts. The layer output is computed via a differential operator, $\mathbf{Z}^{\ell+1}\leftarrow S-\lambda N$, where $\lambda$ is an annealing coefficient that dynamically suppresses the noise subspace, encouraging the model to learn long-term correlation (e.g. causal logic) during training.
@@
$\begin{split}\mathbf{Z}^{\ell+1}\approx&\mathbf{Z}^{\ell}+\eta\underbrace{\text{softmax}\left(\frac{(\mathbf{Z}^{\ell}\mathbf{U}_{s})(\mathbf{Z}^{\ell}\mathbf{U}_{s})^{\top}}{\sqrt{d}}\right)}_{\text{Attention Matrix}}(\mathbf{Z}^{\ell}\mathbf{U}_{s})\mathbf{U}^{\top}_{s}\\ &-\lambda\eta\underbrace{\text{softmax}\left(\frac{(\mathbf{Z}^{\ell}\mathbf{U}_{n})(\mathbf{Z}^{\ell}\mathbf{U}_{n})^{\top}}{\sqrt{d}}\right)}_{\text{Attention Matrix}}(\mathbf{Z}^{\ell}\mathbf{U}_{n})\mathbf{U}_{n}^{\top}.\end{split}$ (6)
Here, the term $\mathbf{Z}^{\ell}\mathbf{U}_{s}$ represents the projection onto the subspace $\mathbf{U}_{s}$, effectively removing the noise if the noise is incoherent with the useful signal.
An illustration of the Prism block is shown in Fig. 1, the input signal is first projected and separated and then modulated by $\pi$-RoPE to capture different features of different scale, and finally processed via differential de-noising to remove the noise.
@@ 2.3 Mean-Field Dynamics and Spectral Decoupling
$f(z)=z+\int\Phi(u,z)\rho_{S}(u)d(u)-\lambda\int\Phi(u,z)\rho_{N}(u)du$ (8)
where $\rho_{S}(u)$ and $\rho_{N}(u)$ are the probability density function of the weights $u$, and the kernel function $\Phi(x,z)$ is the kernel with respect to attention matrix. Although practical models use a finite number of heads, the mean-field formulation provides insights into the asymptotic behavior of the signal/noise separation dynamics.
Consider the mean square error with data $y$,
$\mathcal{L}(\rho_{S},\rho_{N})\approx\frac{1}{2}\mathbb{E}_{z}|y-(\text{Signal}-\lambda\cdot\text{Noise})|^{2},$
under the mean-field limit the Wasserstein gradient flow is given by:
@@
$~~~\left.\left.-\lambda\int K_{SN}d\rho_{N}\right)\right)$
@@
$~~~+\left.\left.\lambda^{2}\int K_{NN}d\rho_{N}\right)\right)$
where $K(u,v)=\int\Phi(u,z)\Phi(v,z)\rho_{z}dz$ is the kernel function characterizing the coherence or similarity between different heads and the $V_{\mathrm{data}}^{S}(u)$ and $V_{\mathrm{data}}^{N}(u)$ are terms with respect to training data, and $K_{SN},K_{NN},K_{SN}$ are the cross-interaction terms between signal and noise heads.
@@
where $\mathbb{E}_{\tau}$ is the integral of different step $\tau$.
@@
$\cos\omega_{u}\tau\cdot\cos\omega_{v}\tau=\frac{1}{2}\cos((\omega_{u}+\omega_{v})\tau)+\cos((\omega_{u}-\omega_{v})\tau)$ (11)
The first term is oscillatory and vanishes as $\tau\xrightarrow{\infty}$ by Riemannian-Lebesgue lemma. Resonance occurs when the frequencies between signal and noise heads align, leading to coupling. To eliminate this, we want the orbit to be dense and uniformly distributed on the torus. Motivated by Weyl’s Equidistribution Theorem  [^2] https://en.wikipedia.org/wiki/Equidistribution_theorem, we suggest choosing the irrational frequency in RoPE. However, mere irrationality may be insufficient; we require not only non-resonance but also that the system is far from resonance. This requirement is analogous to the stability conditions in Hamiltonian mechanics, specifically the Kolmogorov-Arnold-Moser (KAM) theorem  [^3] https://en.wikipedia.org/wiki/Kolmogorov-Arnold-Moser_theorem. In the series expansion of the interaction kernel $K_{SN}$, denominators of the form $\frac{1}{n\omega_{S}-m\omega_{N}}$ appear (where $n,m$ are integers). If this denominator is too small, the system becomes unstable. So we finally suggest using irrational numbers satisfying Diophantine conditions in RoPE  [^4] https://en.wikipedia.org/wiki/Diophantine_approximation The numbers are very hard to approximate by rational numbers.. This leads to our $\pi$-RoPE [^5] One may choose $\phi=\frac{1+\sqrt{5}}{2}$, but the number may be too small to separate the long-range and short-range signal.
@@
where $\theta_{\mathrm{base}}$ is the base frequency [^6] Usually denoted as $b$. in RoPE that is usually chosen to be $10000$ in GPT-2. More theoretical benefit can be shown, if we study the Wasserstein gradient flow of Prism with respect to $\rho_{S}$ and $\rho_{N}$, their evolution equation has an interaction kernel term which can be eliminated by $\pi$-RoPE. Thus, the subspace between signal and noise are nearly decoupled and the whole optimization becomes a good optimization problem. The signal heads find the long-range and low-frequency signal and the noise heads fit the high-frequency, short-range solution.
Finally we adopt $\pi$-RoPE as an ansatz to design our rotary embeddings and give a conjecture below:
Conjecture : (Non-Resonance Condition in Prism). We posit that selecting a rotational frequency ratio $\omega_{S}/\omega_{N}\notin\mathbb{Q}$ (an irrational number, specifically Diophantine Numbers like $\pi$) minimizes the integrated cross-correlation $K_{\text{SN}}$ between the Signal and Noise subspaces over long sequences.

### v2

# PRISM: Deriving a White-Box Transformer as a Signal-Noise Decomposition Operator via Maximum Coding Rate Reduction

@@ Abstract
Deep learning models, particularly Transformers, are often criticized as ”black boxes” and lack interpretability. We propose Prism, a white-box attention-based architecture derived from the principles of Maximizing Coding Rate Reduction ($\text{MCR}^{2}$). By modeling the attention mechanism as a gradient ascent process on a distinct signal-noise manifold, we introduce a specific irrational frequency separation ($\pi$-RoPE) to enforce incoherence between signal (semantic) and noise (syntactic) subspaces. We show empirical evidence that these geometric inductive biases can induce unsupervised functional disentanglement alone. Prism spontaneously specializes its attention heads into spectrally distinct regimes: low-frequency heads capturing long-range causal dependencies (signal) and high-frequency heads handling local syntactic constraints and structural artifacts. To provide a theoretical grounding for these spectral phenomena, we draw an analogy between attention mechanism and a Hamiltonian dynamical system and identify that the standard geometric progression of Rotary Positional Embeddings (RoPE) induces dense resonance networks (Arnold Tongues), leading to feature rank collapse. Empirical validation on 124M-parameter models trained on OpenWebText demonstrates that Prism spontaneously isolates the Attention Sink pathology and maintains isentropic information flow across layers. Further, we suggest a physics-informed plug-and-play intervention KAM-RoPE for large language models (LLMs). Our results suggest that interpretability and performance can be unified through principled geometric construction, offering a theoretically grounded alternative to heuristic architectural modifications
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) has established itself as the fundamental building block of modern artificial intelligence, exhibiting a remarkable power-law relationship between parameter scale and performance (Kaplan et al., 2020). However, despite their empirical success, these models operate largely as black boxes. One limitation is the architectural conflation between semantic reasoning (the ability to manipulate abstract concepts over long-range contexts) and syntactic memorization (the rote statistical completion of local patterns). While recent mechanistic interpretability efforts have identified specific circuits such as induction heads (Olsson et al., 2022), the standard attention mechanism lacks the physical inductive bias required to separate these distinct functional regimes by design.
As these models scale, distinct phenomenological anomalies have emerged. Notably, standard causal Transformers tend to ”dump” excessive probability mass onto the initial token, a phenomenon called ”Attention Sink” (Xiao et al., 2024), and frequently collapse into degenerative repetition loops under long-context generation. We propose that these are not merely engineering artifacts, but symptoms of a feature rank collapse: a structural failure where the attention mechanism, trapped by spectral resonances, loses the capacity to distinguish informative signals from high-frequency noise in deep layers. These limitations may often result in brittle reasoning, hallucination driven by overfitting to noise and data inefficiency.
@@
To enhance interpretability, we turn to first-principles approaches that view deep learning as data compression. The principle of Maximal Coding Rate Reduction ($MCR^{2}$) (Chan et al., 2022) shows that the fundamental goal of representation learning is to expand the volume of the dataset globally while compressing the volume of similar classes locally. Building on this theoretical foundation, the CRATE (Coding-RATE transformer) framework (Yu et al., 2024) establishes that the self-attention mechanism is mathematically isomorphic to a step of gradient ascent optimizing the rate reduction. This perspective transforms the attention layer mechanism from an empirical heuristic into a derived operator for signal denoising on low-dimensional manifolds, providing the theoretical foundation for our architectural interventions.
However, gaps between white-box neural networks and black-box neural networks still exist. One critical gap lies in data modeling; current mathematically interpretable white-box models (e.g., unfolded optimization networks) typically assume clean data with Gaussian noise (Chan et al., 2022), whereas the noise distributions of real-world data are highly non-Gaussian and structural. This discrepancy leads to a conflict in the optimization trajectory. We posit that in the context of global reasoning, the definition of “noise” should be relative, structured and scale-dependent. While local syntactic constraints (e.g., grammar, n-gram statistics) are essential for surface-level linguistic coherence, they manifest as high-frequency interference against the low-frequency global dependencies (semantic signals) required for long-range reasoning. In standard Transformers with limited capacity, the gradient flows of these two regimes undergo spectral entanglement. The high-energy local syntactic artifacts often overshadow the sparse global semantic signals, hindering the optimization process and leading to feature rank collapse.
@@
To mitigate this, we propose Prism, a physics-informed transformer architecture that introduces two specific geometric constraints to the standard Transformer block based on explicit additive noise. First, we impose an overcomplete dictionary (Candès et al., 2006) to expand the representational phase space, allowing signal and noise to separate in higher dimensions without competing for the same limited basis vectors. Second, we implement frequency separation via $\pi$-scaled Rotary Positional Embeddings ($\pi$-RoPE) (Su et al., 2024) to encourage the incoherence between signal and noise based on resonance. By scaling the rotational bases of the overcomplete subspaces by $\pi$ and $1/\pi$ respectively, we propose that this choice effectively mitigates resonance and keeps the reasoning and memorization dynamics on spectrally disjoint manifolds.
@@
For real-world datasets, we validate Prism by scaling it to 124M parameters on the OpenWebText corpus. Despite utilizing standard training recipes, Prism exhibits a distinct thermodynamic signature compared to the GPT-2 baseline. Most notably, we observe the absence of the Attention Sink phenomenon and the preservation of high attention entropy (Isentropic Flow) throughout the network depth, and this can be interpreted by mapping the attention mechanism to a classical Hamiltonian system and studying its perturbation.
Our results suggest that by enforcing rigorous geometric and physical constraints—specifically the orthogonality between signal and noise, we can construct ”white-box” Transformers that are not only interpretable by design but also dynamically robust. Preliminary theoretical analysis suggests a path to physics-informed interventions (KAM-RoPE) for existing LLMs.
@@ 2.1 Rate Reduction as the Objective
and the learning objective is to extract $\mathbf{Z}_{0}$ from $\mathbf{Z}$. To achieve this, we follow the maximum coding rate reduction framework. Let $\mathbf{Z}\in\mathbb{R}^{N\times d}$ denote representation (e.g. input tokens), and $\mathbf{U}_{s},\mathbf{U}_{n}$ be the subspace with respect to the effective signal $\mathbf{Z}_{0}$ and noise $\mathbf{E}$ respectively. We assume the signal and noise reside in distinct subspaces, and these subspaces $\mathbf{U}_{s}$ and $\mathbf{U}_{n}$ are sufficiently incoherent (e.g., $\left\|\mathbf{U}_{s}^{\top}\mathbf{U}_{n}\right\|\approx 0$). We can approximate the noise $\mathbf{E}=\mathbf{Z}\mathbf{U}_{n}$. Thus, the effective signal $Z_{0}$ is obtained by maximizing the coding rate difference $\Delta R(\mathbf{Z})$:
@@
where $R(\mathbf{Z})$ represents the coding rate of the signal projected onto local subspaces expanded by $\mathbf{U}_{s}$, and $R(\mathbf{Z}|\mathbf{U}_{n})$ represents the coding rate of the noise obtained by projecting the noisy signal onto the local subspaces $\mathbf{U}_{n}$, and $\lambda>0$ is the parameter controlling the decomposition between signal and noise. While the closed-form expression of the coding rate is originally derived under the Gaussian assumption, we adopt it here as a robust geometric proxy for measuring the volume of the feature manifold. The coding rate of the matrix $\mathbf{Z}$ is defined as:
@@
If there are $K$ subspaces $\mathbf{U}=[\mathbf{U}_{1},\dots,\mathbf{U}_{K}]$, the coding rate of $\mathbf{Z}$ is the sum of the coding rates of all subspace projections, and yields
$R(\mathbf{Z}\mathbf{U}_{k})=\frac{1}{2}\log\det\left(\mathbf{I}+\beta(\mathbf{Z}\mathbf{U}_{k})(\mathbf{Z}\mathbf{U}_{k})^{\top}\right),$ (4)
where $\beta=\frac{p}{N\epsilon^{2}}$ and $p$ is the dimension of subspaces and $k=1,\dots,K.$
@@ 2.2 Prism as Gradient Ascent of Coding Rate Reduction
Figure 1: The PRISM White-Box Architecture. (Top) The input latent state $\mathbf{Z}^{\ell}$ is projected into an overcomplete feature space via dictionary $\mathbf{U}=[\mathbf{U}_{s},\mathbf{U}_{n}]$. The data flow in Prism bifurcates into Signal Stream (Blue) and Noise Stream (Orange). $\pi$-RoPE: the signal stream applies low-frequency rotary embeddings scaled by $\pi$ ($\theta_{S}=\pi\theta_{\mathrm{base}}$) to represent long-range semantic structures, while the noise stream applies high-frequency embeddings ($\theta_{N}=\frac{1}{\pi}\theta_{\mathrm{base}}$) to capture short-range syntactic artifacts. The ’noise stream’ is explicitly designed to capture and isolate syntactic artifacts or short-range information in NLP tasks. The layer output is computed via a differential operator, $\mathbf{Z}^{\ell+1}\leftarrow S-\lambda N$, where $\lambda$ is an annealing coefficient that dynamically suppresses the noise subspace, encouraging the model to learn long-term correlation (e.g. causal logic) during training.
@@
$\begin{split}\mathbf{Z}^{\ell+1}\approx&\mathbf{Z}^{\ell}+\eta\underbrace{\text{softmax}\left(\frac{(\mathbf{Z}^{\ell}\mathbf{U}_{s})(\mathbf{Z}^{\ell}\mathbf{U}_{s})^{\top}}{\sqrt{d_{k}}}\right)}_{\text{Attention Matrix}}(\mathbf{Z}^{\ell}\mathbf{U}_{s})\mathbf{U}^{\top}_{s}\\ &-\lambda\eta\underbrace{\text{softmax}\left(\frac{(\mathbf{Z}^{\ell}\mathbf{U}_{n})(\mathbf{Z}^{\ell}\mathbf{U}_{n})^{\top}}{\sqrt{d_{k}}}\right)}_{\text{Attention Matrix}}(\mathbf{Z}^{\ell}\mathbf{U}_{n})\mathbf{U}_{n}^{\top}.\end{split}$ (6)
Here, the term $\mathbf{Z}^{\ell}\mathbf{U}_{s}$ represents the projection onto the subspace $\mathbf{U}_{s}$, effectively separating the noise if the noise is incoherent with the useful signal, and $d_{k}$ is the dimension of the space $\mathbf{U}_{s}$ or $\mathbf{U}_{n}$.
An illustration of the Prism block is shown in Fig. 1, the input signal is first projected and separated and then modulated by $\pi$-RoPE to capture different features at different scales, and finally processed via differential de-noising to remove the noise. The incoherence between $U_{s}$ and $U_{n}$ ensures that the denoising process suppresses high-frequency interference without eroding the integrity of the semantic signal accumulated in the residual stream.
@@
While in our derivation, MCR2 formulates representation learning as separating signal from noise, this may not be applied directly for natural language processing (NLP) tasks. We instead reinterpret the signal as semantic subspace corresponding to long-range causal dependency and the noise as syntactic subspace with respect to local constraints, rigid structural artifacts, repetition requiring subtraction. The noise removing process is now reinterpreted as suppressing interference to prevent high-frequency syntactic artifacts (e.g., repetitive loops) from spectrally overshadowing the low-frequency semantic signals. Note that this subtraction operates on the residual update $\Delta Z$, filtering the attention head’s contribution, rather than erasing the syntactic information already accumulated in the residual stream. Under this definition, the ’Attention Sink’ phenomenon is classified as a high-energy syntactic artifact residing in the noise subspace.
@@ 2.3 Mean-Field Dynamics and Spectral Decoupling
$f(z)=z+\int\Phi(u,z)\rho_{S}(u)\,\mathrm{d}u-\lambda\int\Phi(u,z)\rho_{N}(u)\,\mathrm{d}u$
where $\rho_{S}(u)$ and $\rho_{N}(u)$ are the probability density function of the weights $u$, and the kernel function $\Phi(x,z)$ is the kernel with respect to the attention matrix. Although practical models use a finite number of heads, the mean-field formulation provides insights into the asymptotic behavior of the signal/noise separation dynamics.
Considering the mean square error with training data $y,z$,
$\mathcal{L}(\rho_{S},\rho_{N})\approx\frac{1}{2}\mathbb{E}_{z,y}|y-(\text{Signal}-\lambda\cdot\text{Noise})|^{2},$
under the mean-field limit, the Wasserstein gradient flow is given by:
@@
$~~~\left.\left.-\lambda\int K_{SN}\,\mathrm{d}\rho_{N}\right)\right)$
@@
$~~~+\left.\left.\lambda^{2}\int K_{NN}\,\mathrm{d}\rho_{N}\right)\right)$
where $K(u,v)=\int\Phi(u,z)\Phi(v,z)\rho_{z}\,\mathrm{d}z$ is the kernel function characterizing the coherence or similarity between different heads and the $V_{\mathrm{data}}^{S}(u)$ and $V_{\mathrm{data}}^{N}(u)$ are terms with respect to training data, and $K_{SN},K_{NN},K_{NS}$ are the cross-interaction terms between signal and noise heads.
@@
where $\mathbb{E}_{\tau}$ is the integral over different steps $\tau$.
@@
$\cos\omega_{u}\tau\cdot\cos\omega_{v}\tau=\frac{1}{2}\cos((\omega_{u}+\omega_{v})\tau)+\cos((\omega_{u}-\omega_{v})\tau)$
The first term is oscillatory and vanishes as $\tau\xrightarrow{\infty}$ due to phase cancellation (destructive interference) over long contexts, analogous to the Riemann-Lebesgue lemma. Resonance occurs when the frequencies between signal and noise heads align, leading to coupling. To eliminate this, we want the orbit to be dense and uniformly distributed on the torus. Motivated by Weyl’s Equidistribution Theorem (Einsiedler and Ward, 2011), we suggest choosing the irrational frequency in RoPE. However, mere irrationality may be insufficient; we require not only non-resonance but also that the system is far from resonance. This requirement is analogous to the stability conditions in Hamiltonian mechanics, specifically the Kolmogorov-Arnold-Moser (KAM) theorem (Arnold, 1989).
To see this, we note that the cross-interaction kernel $K_{SN}$ and $K_{NS}$ explicitly quantifies the violation of subspace incoherence. While MCR2 requires incoherence ($\mathbf{U}_{S}^{\top}\mathbf{U}_{N}\approx 0$), RoPE can induce subspace alignments dynamically. In the perturbative series analysis aimed at eliminating these interaction terms $K_{SN}$ (via canonical transformation), denominators of the form $\frac{1}{n\omega_{S}-m\omega_{N}}$ appear (where $n,m$ are integers). Standard RoPE frequencies lead to the so-called Small Divisors Problem (resonance), causing the perturbative series to diverge. Physically, this implies that the interaction terms ($K_{SN},K_{NS}$) become non-removable, effectively mixing the signal and noise subspaces over long contexts and violating the MCR2 objective.
To avoid resonance and divergence, we finally suggest using irrational numbers satisfying Diophantine conditions (Einsiedler and Ward, 2011) in RoPE. This leads to our $\pi$-RoPE  [^2] One may choose $\phi=\frac{1+\sqrt{5}}{2}$, but the number may be too small to separate the long-range and short-range signal.
@@
where $\theta_{\mathrm{base}}$ is the base frequency  [^3] Usually denoted as $b$ with typical value $10000$. in RoPE that is usually chosen to be $10000$ in GPT-2. More theoretical benefit can be shown, if we study the Wasserstein gradient flow of Prism with respect to $\rho_{S}$ and $\rho_{N}$, their evolution equation has an interaction kernel term which can be eliminated by $\pi$-RoPE. Thus, the subspaces between signal and noise are nearly decoupled and the whole optimization becomes a good optimization problem. The signal heads find the long-range and low-frequency signal and the noise heads fit the high-frequency, short-range solution.
Finally, we adopt $\pi$-RoPE as an ansatz to design our rotary embeddings and give a conjecture below:
Conjecture: (Non-Resonance Condition in Prism). We posit that selecting a rotational frequency ratio $\omega_{S}/\omega_{N}\notin\mathbb{Q}$ (an irrational number, specifically irrational numbers satisfying Diophantine conditions such as $\pi$ or $\phi$) minimizes the integrated cross-correlation $K_{\text{SN}},K_{\text{NS}}$ between the signal and noise subspaces over long sequences.